OpenStudy (anonymous):

The slope of a line through (15, -8) that is perpendicular to the line with the equation y= 5x +9

OpenStudy (praxer):

if the lines are perp. then . slope of line 1 * slope of line 2 = -1 so, slope of line 1= -1/slope of line 2 = -1/5 now the equation of line 1 is y-y1= slope * (x-x1) => y-(-8)=-1/5*(x-16) => y+8=(16-x)/5 =>5y+x-24=0 the desired equation is $$\huge \color {red}{5y+x-24=0}$$

OpenStudy (praxer):

sorry, I have taken x1 as 16. Do substitute it and do the calculations :P

OpenStudy (anonymous):

If the slope of the line is -1/5 what is the y intercept

OpenStudy (praxer):

for the y intercept in the desired equation put x= 0 the value for y will give the y intercept ...

OpenStudy (praxer):

after plugging x1 as 15 u will get, $$\huge \color{red}{5y+x-25=0}$$